# Exact treatment of the quantum Langevin equation under time-dependent system-bath coupling via a train of delta distributions

### v1

@@ Abstract
In this paper, we consider the quantum Langevin equation for the Caldeira-Leggett model with an arbitrary time-dependent coupling constant. We solve this equation exactly by employing a train of Dirac-delta switchings. This method also enables us to visualize the memory effect in the environment. Furthermore, we compute the two-time correlation functions of the system’s quadratures and show that the discrete-time Fourier transform is well-suited for defining spectral densities, as the Dirac-delta switchings turn continuous functions into discretized samples.
@@ I Introduction
The Caldeira-Leggett model [4, 5, 6] is a cornerstone in the theory of open quantum systems, describing the dynamics of a quantum Brownian particle [2]. In this framework, a quantum harmonic oscillator serves as the system, and it is coupled to an environment composed of a collection of quantum harmonic oscillators. Such a model is extremely useful for analyzing the behavior of open quantum systems because it allows for an exact solution to the equation of motion, known as the quantum Langevin equation (QLE) [1, 7, 8].
@@
In many studies, the focus is on the asymptotic behavior of the system, e.g., the two-time correlation functions of the system quadratures in the long-interaction limit, during which the system reaches equilibrium with the environment. Therefore, it is assumed that the system constantly couples to the bath for an infinitely long time. In fact, this time-independent coupling is crucial for solving the QLE analytically. However, there have been few studies on the QLE with time-dependent coupling (i.e., non-stationary QLE). The primary reason is that obtaining an analytical solution is challenging. While some research applies Floquet theory to study the asymptotic behavior of periodically driven systems obeying the QLE [9, 10, 11, 12], the solution to the QLE with an arbitrary time-dependent coupling is still unknown. [^1] We note that, in the field of coarse-grained molecular dynamics simulation, the generalized Langevin equation (GLE) in the Mori-Zwanzig theory has been extensively studied. In recent years, non-stationary GLE has been analyzed by using the so-called memory reconstruction method [13, 14, 15]. This issue is particularly relevant in the study of finite-time interactions in, e.g., quantum thermodynamics [16].

### v3

# Exact treatment of the memory kernel under time-dependent system-environment coupling via a train of delta distributions

@@ Abstract
Memory effects in a quantum system coupled to an environment are one of the central features in the theory of open quantum systems. The dynamics of such quantum systems are typically governed by an equation of motion with a time-convolution integral of the memory kernel. However, solving such integro-differential equations is challenging, especially when the memory kernel is nonstationary (not time-translation invariant). In this paper, we analytically and nonperturbatively solve such integro-differential equations with a nonstationary memory kernel by employing a train of Dirac-delta switchings. We then apply this method to the damped Jaynes-Cummings model and the damped harmonic oscillator model to demonstrate that (i) our solution asymptotes to the well-known exact solution in the continuum limit, and that (ii) our method also enables us to visualize the memory effect in the environment.
@@ I Introduction
Mathematically, such memory effects appear in equations of motion and cause difficulties. We are particularly interested in a quantity called the memory kernel $\Sigma(t,t^{\prime})$, which appears as a time-convolution integral (“an integral over the past history of the system” [5]):
$\int_{t_{0}}^{t}\text{d}t^{\prime}\,\Sigma(t,t^{\prime})\mathcal{O}(t^{\prime})\,$, (1)
where $\mathcal{O}(t)$ is some observable. This structure arises, for example, in the Nakajima-Zwanzig equation [23, 32]—an exact integro-differential equation for the density matrix of the system of interest. Because such integro-differential equations are difficult to solve in general, analyses often rely on perturbation theory.
There are, however, special models that admit exact solutions even in the presence of a time-convolution integral. One example is a two-level system immersed in a bosonic bath. In particular, the damped Jaynes–Cummings model describes the coupling of a two-level system to a single-mode bosonic bath with a Lorentz spectral density. Another example is a damped harmonic oscillator obeying the Heisenberg equation of motion known as the quantum Langevin equation (QLE) [30, 15, 16]. Specifically, the
@@
Caldeira–Leggett model [6, 7, 8] is a cornerstone in the theory of open quantum systems, describing the dynamics of a quantum Brownian particle [5]. In this framework, a quantum harmonic oscillator serves as the system, and it is coupled to an environment composed of a collection of quantum harmonic oscillators.
@@
While these exactly solvable models are extremely useful, they rely on the stationarity of the memory kernel. In many studies, the focus is on the asymptotic behavior of the system, e.g., the density matrix or the two-time correlation functions of the system quadratures in the long-interaction limit, during which the system reaches equilibrium with the environment. Therefore, it is assumed that the system constantly couples to the bath for an infinitely long time. In fact, this time-independent coupling is crucial for solving the master equation and QLE analytically, as the memory kernel in these models is time-translation invariant (i.e., stationary), $\Sigma(t,t^{\prime})=\Sigma(t-t^{\prime})$, and so the time-convolution integral (1) can be treated without approximations. However, once we introduce a time-dependent coupling, the memory kernel becomes nonstationary, which makes it challenging to solve the equation of motion analytically. This issue is particularly relevant in the study of finite-time interactions in, e.g., quantum thermodynamics [9].
In this paper, we propose a nonperturbative method for solving equations of motion with the time-convolution integral, even when the memory kernel is nonstationary. In particular, we consider a class of integro-differential equations that contains the memory kernel of the form
$\Sigma(t,t^{\prime})=\chi(t)\chi(t^{\prime})\Gamma(t,t^{\prime})\,$, (2)